# Gaudin models and moduli space of flower curves

### v1

@@ 7.2. Quadratic parts of inhomogeneous/trigonometric Gaudin subalgebras
Theorem 8.2 ([R4]).
Suppose that $C\in\overline{M}_{n+1}^{split}$. Then the subalgebra $\mathcal{A}(C)$ acts by normal operators, hence with simple spectrum, on $V({\underline{\lambda}})$.
@@
We need a generalization of Proposition 8.2. Let $\theta\in\mathfrak{h}^{split}$. Consider the Hermitian version of the Shapovalov form on $M(\theta)$, i.e. the Hermitian form such that for the highest vector $v_{\theta}\in M(\theta)$ we have
@@
Note that the Shapovalov form for finite-dimensional representation coincides with the standard Hermitian form. Define the Hermitian form on $M(\theta)\otimes V(\lambda_{1})\otimes\ldots V(\lambda_{n-1})$ as the product of the above Shapovalov form on $M(\theta)$ and standard Hermitian forms on $V(\lambda_{i}),i=1,\ldots,n-1$.
@@
The restriction of the above Hermitian form to $(M(\theta)\otimes V(\lambda_{1})\otimes\ldots V(\lambda_{n-1}))^{\mathfrak{n}_{+}}$ for big enough $\theta\in\mathfrak{h}^{split}$ is positive definite.
@@
For any $\theta$, the tensor product $M(\theta)\otimes V(\lambda_{1})\otimes\ldots V(\lambda_{n-1})$ is is an object from the category $\mathcal{O}$ which has a standard filtration with the subquotients being the Verma modules with the highest weights $\theta+\mu$, for all weights $\mu$ of $V(\lambda_{1})\otimes\ldots V(\lambda_{n-1})$. It follows that any singular vector belongs to a finite sum of weight spaces. On each weight space, the above form degenerates for $\theta$ from a finite union of hyperplanes parallel to the root hyperplanes [S]. On the other hand, for integral dominant $\theta$ such that for any positive root $\alpha$, the value of $(\theta,\alpha)$ is big enough, the above Hermitian form coincides with the standard one on the weight space in the tensor product of finite dimensional $\mathfrak{g}$-modules, so is positive definite. This means that for any $\theta\in\mathfrak{h}^{split}$ such that for any positive root $\alpha$, the value of $(\theta,\alpha)$ is big enough, this form is positive definite as well. ∎
@@
For any $C\in\overline{M}_{n+1}^{split}$ and big dominant $\theta\in\mathfrak{h}^{split}$, the subalgebra $\mathcal{A}(C)$ acts by normal operators on $(M(\theta)\otimes V(\lambda_{1})\otimes\ldots V(\lambda_{n-1}))^{\mathfrak{n}_{+}}$.
@@
Note that the condition on operators to be normal is a closed condition hence it is enough to check the statement for $\mathcal{A}(0,z_{1},\ldots,z_{n})$ with $z_{1},\ldots,z_{n}\in\mathbb{R}$. From the proposition above we have a positive defined Hermitian form on $(M(\theta)\otimes V(\lambda_{1})\otimes\ldots V(\lambda_{n-1}))^{\mathfrak{n}_{+}}$. From Proposition 4.4, we know that the universal Gaudin subalgebra $\mathcal{A}_{\mathfrak{g}}$ is invariant under the anti-involution $f_{i}[k]\mapsto e_{i}[k]$ and $h_{i}[k]\to h_{i}[k]$, so for any $x\in\mathcal{A}(C)$ we have $x^{+}\in\mathcal{A}(C)$, i.e. all the operators coming from $\mathcal{A}(C)$ are normal with respect to the Hermitian form. ∎
@@
Note that if $\theta=\varepsilon^{-1}\chi$ then $\mathcal{A}_{\theta}^{trig}(C)=\mathbfcal{A}^{\varepsilon}_{\chi}(C)$, so it suffices to prove the statement for $\mathbfcal{A}^{\varepsilon}_{\chi}(C)$. Let $Y$ be the subset of ¯$\overline{\mathcal{F}}_{n}$ consisting of such $(C;\varepsilon)$ that there is no cyclic vector in $V({\underline{\lambda}})$ for the corresponding subalgebra $\mathbfcal{A}^{\varepsilon}_{\chi}(C)$. This set is Zariski closed. The map
@@
The family of trigonometric Gaudin subalgebras $\mathcal{A}_{\theta}^{trig}(C)$ with $C\in\overline{M}_{n+2}^{split}$ and big enough generic dominant $\theta\in\mathbb{R}\chi\subset(\mathfrak{h}^{split})^{reg}$ acts with simple spectrum on $V({\underline{\lambda}})$.
@@
    Identifying the Hom-space $\text{Hom}_{\mathfrak{g}}((M(\theta+\mu),M(\theta)\otimes V({\underline{\lambda}}))$ with the coinvariants space $(M(\theta+\mu)^{*}\otimes V({\underline{\lambda}}))_{\mathfrak{n}_{+}}^{-\theta}$. The image of $\mathcal{A}(u_{\infty},u_{0},u_{1},\ldots,u_{n})$ is then the image of the Gaudin algebra $\mathcal{A}(0,z_{1}^{-1},\ldots,z_{n}^{-1})$ which factors through $\mathcal{A}^{trig}_{-\theta-\mu}(z_{1}^{-1},\ldots,z_{n}^{-1})$, since $-\theta-\mu$ is the lowest weight of $M(\theta+\mu)^{*}$, so it is $\pi_{\underline{\lambda}}^{\mu}(\mathcal{A}^{trig}_{-\theta-\mu}(z_{1}^{-1},\ldots,z_{n}^{-1}))$.

### v2

@@ 7.2. Quadratic parts of inhomogeneous/trigonometric Gaudin subalgebras
Theorem 8.2 ([FFRy]).
Suppose that $C\in\overline{M}_{n+1}^{split}$. Then the subalgebra $\mathcal{A}(C)$ acts by normal operators on $V({\underline{\lambda}})$.
@@
We will need a generalization of Theorem 8.2. Let $\theta\in\mathfrak{h}^{split}$. Consider the Hermitian version of the Shapovalov form on $M(\theta)$, i.e. the Hermitian form such that for the highest vector $v_{\theta}\in M(\theta)$ we have
@@
Note that the Shapovalov form for finite-dimensional representation coincides with the standard Hermitian form. Define the Hermitian form on $M(\theta)\otimes V(\lambda_{1})\otimes\ldots\otimes V(\lambda_{n-1})$ as the product of the above Shapovalov form on $M(\theta)$ and standard Hermitian forms on $V(\lambda_{i}),i=1,\ldots,n-1$.
@@
The restriction of the above Hermitian form to $(M(\theta)\otimes V(\lambda_{1})\otimes\ldots\otimes V(\lambda_{n-1}))^{\mathfrak{n}_{+}}$ for big enough $\theta\in\mathfrak{h}^{split}$ is positive definite.
@@
For any $\theta$, the tensor product $M(\theta)\otimes V(\lambda_{1})\otimes\ldots\otimes V(\lambda_{n-1})$ is is an object from the category $\mathcal{O}$ which has a standard filtration with the subquotients being the Verma modules with the highest weights $\theta+\mu$, for all weights $\mu$ of $V(\lambda_{1})\otimes\ldots V(\lambda_{n-1})$. It follows that any singular vector belongs to a finite sum of weight spaces. On each weight space, the above form degenerates for $\theta$ from a finite union of hyperplanes parallel to the root hyperplanes [S]. On the other hand, for integral dominant $\theta$ such that for any positive root $\alpha$, the value of $(\theta,\alpha)$ is big enough, the above Hermitian form coincides with the standard one on the weight space in the tensor product of finite dimensional $\mathfrak{g}$-modules, so is positive definite. This means that for any $\theta\in\mathfrak{h}^{split}$ such that for any positive root $\alpha$, the value of $(\theta,\alpha)$ is big enough, this form is positive definite as well. ∎
@@
For any $C\in\overline{M}_{n+1}^{split}$ and big dominant $\theta\in\mathfrak{h}^{split}$, the subalgebra $\mathcal{A}(C)$ acts by normal operators on $(M(\theta)\otimes V(\lambda_{1})\otimes\ldots\otimes V(\lambda_{n-1}))^{\mathfrak{n}_{+}}$.
@@
Note that the condition on operators to be normal is a closed condition hence it is enough to check the statement for $\mathcal{A}(0,z_{1},\ldots,z_{n})$ with $z_{1},\ldots,z_{n}\in\mathbb{R}$. From the proposition above we have a positive defined Hermitian form on $(M(\theta)\otimes V(\lambda_{1})\otimes\ldots\otimes V(\lambda_{n-1}))^{\mathfrak{n}_{+}}$. From Proposition 4.4, we know that the universal Gaudin subalgebra $\mathcal{A}_{\mathfrak{g}}$ is invariant under the anti-involution $f_{i}[k]\mapsto e_{i}[k]$ and $h_{i}[k]\to h_{i}[k]$, so for any $x\in\mathcal{A}(C)$ we have $x^{+}\in\mathcal{A}(C)$, i.e. all the operators coming from $\mathcal{A}(C)$ are normal with respect to the Hermitian form. ∎
@@
Note that if $\theta=\varepsilon^{-1}\chi$ then $\mathcal{A}_{\theta}^{trig}(C)=\mathbfcal{A}^{\varepsilon}_{\chi}(C)$, so it suffices to prove the statement for $\mathbfcal{A}^{\varepsilon}_{\chi}(C)$. Let $Y$ be the subset of $\overline{\mathcal{F}}_{n}$ consisting of such $(C;\varepsilon)$ that there is no cyclic vector in $V({\underline{\lambda}})$ for the corresponding subalgebra $\mathbfcal{A}^{\varepsilon}_{\chi}(C)$. This set is Zariski closed. The map
@@
The family of trigonometric Gaudin subalgebras $\mathcal{A}_{\theta}^{trig}(C)$ with $C\in\overline{M}_{n+2}^{split}$ and big enough $(\theta>>0)$ generic dominant $\theta\in\mathbb{R}\chi\subset(\mathfrak{h}^{split})^{reg}$ acts with simple spectrum on $V({\underline{\lambda}})$.
@@
    Identifying the Hom-space $\text{Hom}_{\mathfrak{g}}((M(\theta+\mu),M(\theta)\otimes V({\underline{\lambda}}))$ with the coinvariants space $(M(\theta+\mu)^{*}\otimes V({\underline{\lambda}}))_{\mathfrak{n}_{+}}^{-\theta}$. The image of $\mathcal{A}(u_{\infty},u_{0},u_{1},\ldots,u_{n})$ is then the image of the Gaudin algebra $\mathcal{A}(0,z_{1}^{-1},\ldots,z_{n}^{-1})$ which factors through $\mathcal{A}^{trig}_{-\theta-\mu}(z_{1}^{-1},\ldots,z_{n}^{-1})$, since $\theta-\mu$ is the lowest weight of $M(\theta+\mu)^{*}$, so it is $\pi_{\underline{\lambda}}^{\mu}(\mathcal{A}^{trig}_{-\theta-\mu}(z_{1}^{-1},\ldots,z_{n}^{-1}))$.